Grade 12Physical Chemistry

Copper sulphate solution is electrolysed for 20 minutes with a current of 3 amperes. What mass of copper will be deposited at the cathode? (EQ. mass of copper=31.75)

SWETA TRIPATHY
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
I = 3 Ampere  
Time = 20 x 60s = 1200s  
Q = I x t  = 3x 1200 = 3600 C  
Cu2+ + 2e- --> Cu (s)
 
hence copper deposited = 63.5 * 3600/2 * 96500 = 228600/(193000) = 1.184 gm